if the total cost of the firm is given by TC=Q2+20
A. The average total coast (ATC)
B. The marginal cost (MC)
C. The average variable cost (AVC)
D. The average fixed cost(AFC)
1
Expert's answer
2019-12-16T09:54:16-0500

TC=Q2+20

A. The average total coast (ATC) = ТС / Q = (Q2+20) / Q

B. The marginal cost (MC) = 2Q (MC is the first derivative of TC with respect to Q)

C. The average variable cost (AVC) = TVC / Q = Q2 / Q = Q

D. The average fixed cost(AFC) = TFC / Q = 20 / Q